OpenStudy (pinklion23):

8th grade math help pls

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

5x+y=5

OpenStudy (pinklion23):

?

OpenStudy (pinklion23):

@imqwerty

OpenStudy (anonymous):

u have to substitute one of them for x or y

OpenStudy (pinklion23):

i know that but i don't have time to go through the whole number grid

OpenStudy (anonymous):

so the second one is good bc there is no number

OpenStudy (pinklion23):

what are you taking about :(

OpenStudy (pinklion23):

DX

OpenStudy (anonymous):

it is xd

OpenStudy (pinklion23):

i don't think you guys understand the question i asked

OpenStudy (anonymous):

x=y+3 so u plug it in the other one 4(y+3)+y=2 4y+12=2 4y=-10 y=-10/4 y=-5/2

imqwerty (imqwerty):

u have these 2 equations-\[4x+y=2 \]\[x-y=3\] add them together\[4x+y+x-y=2+3\]\[5x=5\]\[\frac{ 5x }{5 }=\frac{ 5 }{ 5 }\]\[x=1\]now we have x=1 put this value of x in any of the original equations to get y like i choose to put it in the second equation so\[x-y=3\]\[1-y=3\]\[y=-2\]

OpenStudy (anonymous):

no

OpenStudy (anonymous):

1,-2

OpenStudy (pinklion23):

thx imqwerty

OpenStudy (anonymous):

so this is =\[\frac{ 5 }{ 2 }\]